OpenStudy (anonymous):

Hello everyone, I am suppose to identify the slope, y-intercept, and then graph the line, all they gave me was y=4x, anyone have a clue how to solve this?

OpenStudy (anonymous):

If the equation is the line is \[\Large y=mx+b\] where : m is the slope and b is the y-intercept.

OpenStudy (shamim):

u hv to compare it with the equation\[y=mx+c\]

OpenStudy (anonymous):

so the slope is 4?

OpenStudy (anonymous):

@violetanna Of course !

OpenStudy (anonymous):

and the y intercept is 0?

OpenStudy (anonymous):

@violetanna exactly

OpenStudy (anonymous):

so the graph would look like 

OpenStudy (anonymous):

cuz slope= rise/run so it would be 4/1 and I would just go up 4 and over one when graphing the line correct?

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

You can graph the line in another way by giving two values of x and finding the corresponding values of y like this : \[\Large\text{The equation of the line is : }y=4x.\\ \\\Large \text{ when }x=0\text{ then } y=4\times 0=0 \\\Large\text{ when } x=1 \text{ then } y=4\times 1=4 \] So we place the points (0,0) and (1,4) and we trace the line passing these points

OpenStudy (anonymous):

but can I graph it as slope=rise/run=4/1

OpenStudy (anonymous):

yes And what did you write above is correct. I just gave you another way ;)

OpenStudy (anonymous):

ty